The lengths of the sides of a triangle are consecutive odd integers.
if the perimeter is one less than four times the shortest side. find the length of each side
***please write the equation used to solve this**

Answer by addingup(3677)

Consecutive odd integers. How many numbers from 1 to3? From 3 to 5? From 5 to 7? The answer is two, there are two numbers between odds just like there are two numbers between evens: one, two AND three; three four ANd five.
So we have that:
a+a+2+a+2+2= 4a-1
3a+6= 4a-1 Add 1 to both sides:
3a+7= 4a Subtract 3a from both sides
7= 4a
or
4a= 7 Now divide both sides by 4
a= 7/4 The denominator is bigger than the numerator, it means we have a mixed number:
4 goes into 7 one time and we have a remainder of 3:
1 3/4 Since we are talking length, let's convert to decimal: 3/4= 0.75:
1.75 is the length of the shortest side, let's call it side a
side b: 1.75+2= 3.75
side c: 1.75+2+2= 5.75